OpenStudy (anonymous):

$6500 invested at 6% compounded continuously for 2 years. A=6500e^.12 then what?

terenzreignz (terenzreignz):

You just evaluate it on a calculator, I suppose..

OpenStudy (anonymous):

i cant use a calculator for this class

OpenStudy (mindblast3r):

I don't understand. What are you having trouble with exactly?

terenzreignz (terenzreignz):

That's just sadistic D: No (normal) human can evaluate e^0.12 without the help of technology (right?)

OpenStudy (mindblast3r):

I agree with terenzreignsz

OpenStudy (anonymous):

im supposed to use the natural log and get as close as possible

OpenStudy (anonymous):

a woman invests $6500 into an accoun that pays 6% interest per year, compounded continuously what is the amount after 2 years?

OpenStudy (mindblast3r):

I don't know how to do this without a calculator, sorry.

OpenStudy (anonymous):

so using A=Pe^rt i have A=6500e^(.06)(2)

OpenStudy (anonymous):

what does the A stand for in this formula?

OpenStudy (mindblast3r):

Amount

OpenStudy (mindblast3r):

What country do you live in, if I may ask?

OpenStudy (mindblast3r):

Is E the currency?

OpenStudy (anonymous):

so it is in fact my unknown? USA

OpenStudy (anonymous):

e is part of the formula

OpenStudy (wolf1728):

I think I can help you out

OpenStudy (anonymous):

so the next part of the problem is how long will it take for the amount to be $8000 so i would have 8000=6500e^(.06)(t) then 8000/6500=e^(.06)(t) then 16/13=e^(.06)(t) then ln(16/13)=(.06)(t)

terenzreignz (terenzreignz):

And then divide both sides by 0.06, and that's done :D

OpenStudy (anonymous):

k so the first part that i was originally asking about is impossible to solve with out a calculator?

OpenStudy (agent0smith):

if you took ln of both sides A=6500e^(.12) \[\Large \ln A = \ln (6500e^{0.12}) = \ln 6500 + 0.12 \ln e\]which you can probably solve for A using a table of logs or something. But i'd probably seriously consider getting the whole class to agree on telling the professor to F off with this no-calculator bs.

OpenStudy (anonymous):

its not just the professor its university wide, math 1050 is a no calculator class, it sucks but its how it is..

OpenStudy (wolf1728):

You could solve continuously compounded interest with a logarithm table. I'd help you out some more, but I'm researching this.

OpenStudy (anonymous):

thanks wolf.

OpenStudy (agent0smith):

That seems like a colossal waste of student's time and effort, like forcing students to do long division (other than polynomial) anytime after about 8th grade. Yay for wasted time and energy.

OpenStudy (anonymous):

yeah ive been doing this one assignment since 10:00 pm its now 2:15 am and im still not finished

OpenStudy (agent0smith):

I looked up a couple of math 1050 courses and a couple don't allow graphing calcs, but allow standard scientific ones. And it seems like the ones that didn't allow calculators had questions that didn't require them (like this one. it does require one). You're sure it's not just graphing calcs that aren't allowed?

OpenStudy (anonymous):

im sure, its no calculators at all trust me i wish i could use a calculator

OpenStudy (anonymous):

maybe my professor only wants me to do part b for this problem cuz i can isolate "t" and get my simplified answer

OpenStudy (anonymous):

but idk how he wants me to do part a without a calculator.

OpenStudy (wolf1728):

kylhea - could you read the private message I just sent you?

OpenStudy (anonymous):

thanks wolf

OpenStudy (wolf1728):

Okay kylhea how much longer will you be online here?

OpenStudy (wolf1728):

(Just wondering because I need a little time for the research and the calculation.)

OpenStudy (wolf1728):

Guess I'll start the calculations.

OpenStudy (anonymous):

ill be on a while i need to finish this assignment

OpenStudy (anonymous):

well actually im gonna need to finish it tomorrow, i gotta be up in 3 hours for work

OpenStudy (wolf1728):

Okay - check back here tomorrow - or maybe I can send it in a private message. Have a good night :-)

OpenStudy (anonymous):

ok thanks a ton :) have a great night

OpenStudy (wolf1728):

you too

OpenStudy (wolf1728):

Okay I came up with this 6,500 Principal 6.00% Compounded Continuously 2 years Formula for Converting Annual Rate to Continuously Compounded Rate Continuously compounded rate = (e^r) -1 where e = 2.7182818285 ('e' is the base of natural logarithms). Continuously compounded rate = (2.7182818285^.06) -1 Continuously compounded rate = 1.0618365465 -1 Continuously compounded rate = 0.0618365465 (or 6.18365465% as it is reported in newspapers) Formula for calculating the total amount for an initial amount (principal) invested at a compounded interest rate: Total = Principal * (1 +r)^years Total = 6,500 * (1.0618365465)^2 Total = 6,500 * 1.1274968516 Total = 7,328.73 and yes it's just that simple LOL

OpenStudy (agent0smith):

Continuously compounded rate = (2.7182818285^.06) -1 Continuously compounded rate = 1.0618365465 -1 how'd you do this w/o a calculator?

OpenStudy (wolf1728):

No I did use a calculator but it can be done with a logarithm table.

OpenStudy (wolf1728):

Here's one way to calculate continuously compounded interest - but it talks about generating approximate logs and without using without a log table. (Basically you create your own log table.) http://answers.yahoo.com/question/index?qid=1006051509017 Here's a link to a common logarithm table: http://www.sosmath.com/tables/logtable/logtable.html As agentsmith was so anxious to know, here's how to convert a 6% annual rate to a compounded rate: continuously compounded rate = (e^annual rate) -1 We will have to raise 'e' to the power of .06, then subtract 1. (2.7182818285^.06) -1 = .06 * log (2.72) (table just has 3 sig figures) =.06 * 0.4346 =0.026076 now we look up the ANTI-LOG of .026076 (We look at the bold numbers and read to the left) we see that the answer lies between 1.061 0.02571538 and 1.062 0.02612452 to get the next DECIMAL we look at the logarithm difference = 0.02612452 -0.02571538 = 0.00040914 Meaning that for every .0001LOG increase, the ANTI-LOG increases by 0.000040914 (Notice we have decreased this number by 10) (Yes using Excel here because this is getting WAY too tedious) Log Anti-Log 1.0610 0.025715380 1.0611 0.025756294 1.0612 0.025797208 1.0613 0.025838122 1.0614 0.025879036 1.0615 0.025919950 1.0616 0.025960864 1.0617 0.026001778 1.0618 0.026042692 1.0619 0.026083606 1.0620 0.026124520 Remember we have to find the ANTI-LOG of 0.026076 Now we see that it lies between 1.0618 and 1.0619 (We should subtract 1 to get .0618 AND .0619) So, for just that teensy bit of work, we have narrowed down the continuously compounded to 3 significant figures. Wow, what amazing accuracy.

OpenStudy (wolf1728):

Gee I almost forgot there was a Part B to this problem. $6,500 invested at 6% compounded continuously for 2 years how long will it take for the amount to be $8000? (The 6% interest is (if you remember) is 6.18365465% compounded continuously.) The compound Interest Formula is: Total = Principal × (1 + Rate)^years Solving this for years Years = {log(total) -log(Principal)} ÷ log(1 + rate) Years = {log (8,000) -log(6,500)} ÷ log (1.0618365465) Years = {3.903089987 - 3.8129133566) ÷ 0.0260576689 Years = 0.090176630 ÷ 0.0260576689 Years = 3.460656082 Formulas Used: http://1728.org/compint2.htm Calculator to check answer: http://1728.org/compint.htm